# Compactifications of phylogenetic systems and electrical networks

### v3

@@ Abstract.
We describe a graphical map: a visual and numerical correspondence between circular electrical networks and circular planar split systems. When restricted to the planar circular electrical case, this graphical map finds the split system uniquely associated with the Kalmanson resistance distance of the dual network, matching the induced split system familiar from phylogenetics. This correspondence is extended to compactifications of the respective spaces, taking cactus networks to the cactus split systems defined herein. The graphical map preserves both network components and cactus structure, allowing an elegant enumeration of induced phylogenetic split systems. We introduce the global spaces of circular planar electrical networks and circular split systems. These are CW complexes, and the 0-cells of each are counted by the Bell numbers as opposed to the Catalan numbers.
@@ 1.1. Motivation
Recently it was demonstrated that a favorite tool of phylogenetics researchers, the split system, is guaranteed to exhibit the large scale structure of a simple electrical resistor network [13]. This highlights a mathematical analogy between two problems: reconstructing a circuit in a ‘black box’ from surface measurements, and rebuilding a genetic history from existing population data.
@@
Figure 1. Example labels for the cells in the spaces we consider, for $n=4$: (a) circular planar electrical networks, (b) cactus networks (c) circular split systems and (d) cactus circular split systems . The total numbers of these cells are listed in Table 1, row 4, and the entire sets for (a) and (c) are shown in Figure 4.
@@
In this paper we carry the analogy further. A printed electrical circuit is made of conductors in a network, with terminal nodes on the exterior available for connections. Some of those exterior nodes might be duplicates, short-circuited together so that they are interchangeable in terms of their use. Similarly, monozygotic twins can be seen as genetically the same for the purpose of reconstructing a family tree from the DNA samples of living individuals. On a larger scale, basic DNA sequences of extant individuals from a biologically diverse ecosystem can be equated when the species match. Allowing nodes, either terminals or taxa, to be identical in the limit of being infinitely close to each other corresponds to taking an unbounded space and compactifying it. Here we define and relate the compactification of phylogenetic split systems (from the DNA examples) to the well-studied compactification of circular planar electrical network space.
1 1 1 1 1 1 1
2 2 3 2 3 2 3
3 8 15 8 15 8 15
4 52 105 64 117 49 102
5 464 945 1024 1565 373 839
OEIS [A111088] [A001147] [A006125] [A136654] [?] [?]
Table 1. Spaces: numbers of cells in each CW complex. Left to right these are circular planar electrical networks, cactus planar electrical networks, circular split systems, and cactus circular split systems. The last two columns count the cells in the range of our maps from electrical networks to split systems. Examples are in Figure 1.
From a geometric combinatorial perspective, our interest lies in comparing the spaces of electrical networks and split systems as CW-complexes with analogous cell-decompositions. We will review the definitions in the following sections, but for now we show some examples of each, to introduce our notation. $\mathbf{\Omega}_{n}$ is the space of circular planar electrical networks with $n$ exterior nodes, in notation from [7]. $\mathbf{\Omega}_{n}$ has its poset of cells denoted $EP_{n}$ in [1]. The cells correspond to equivalence classes of critical networks such as (a) in Figure 1. Subfigure 1 (c) shows an example of a split system in $\mathbf{\Psi}_{4}.$ The compactification of $\mathbf{\Omega}_{n}$ is the space of cactus networks, denoted $E_{n}$ in [33]. The cells of $E_{n}$ we denote as $\mathbf{\Omega}_{n},$ with an example in (b) of Figure 1. The compactification of $\mathbf{\Psi}_{n}$ is defined in this paper and is denoted $\mathbf{\Psi}_{n}$, with an example in (d). In the interest of readability, we often use one name for both the CW-complex and the poset of cells: we will refer to $|\mathbf{\Omega}_{n}|$ as the number of cells in $\mathbf{\Omega}_{n}$ and to $|\mathbf{\Psi}_{n}|$ as the number of cells in $\mathbf{\Psi}_{n}.$ Similarly, we will use $\mathbf{\Omega}_{n}$ interchangeably with $E_{n}.$
Further, we want to consider the sub-complexes of split systems that are the range of the graphical map $\mathbf{\xi}$ which we describe and relate to the Kalmanson and induced phylogenetic mappings. In Figure 1 (c) is the image of (a) under $\mathbf{\xi}$, and (d) is the image of (b). There are fewer cells in that range than split systems in general: the implied mapping on cells is not onto. Neither is it an injection of cells. However the graphical map does preserve the connected components and cactus structures, and this provides us a way to count cells in the range either in either case, whether we want just the ordinary split systems or their compactifications as well. Table 1 shows the results of counting the cells of all dimension in the various complexes.
In our recent papers [13, 14], we demonstrated an injective mapping from electrical equivalence classes of circular planar electrical networks to circular split systems. Every planar response matrix is a Kalmanson metric, and thus a unique weighted circular split system. To see if a given circular split system is the image of an electrical network is more complicated. For a given Kalmanson metric one can first reorder the matrix to match any cyclic order respected by its circular split system. Then there is a unique associated response matrix, which can be checked for nonnegative circular minors to decide if it is a planar electrical network. If so, then we show in [13] that the circular planar electrical network has the same overall tree structure as the split system network of the resistance matrix. Then the algorithm from Chapter 9 of [7] can reconstruct a critical subnetwork for any part not made of bridges, using the strand diagram. In [24] the authors show how to check for planarity directly using the (reordered) resistance metric. This is done via the Lam embedding into the nonnegative Grassmannian, but with a variation that takes the resistance metric as input. Then in [24] the authors show an alternative reconstruction of the electrical network again using the resistance metric directly to find the strand matching.
Our map $\mathbf{\xi}$ takes circular planar electrical cactus networks to cactus circular split systems.The map is quite simple in its operation on matrices: it is the identity map. This is true for the cactus networks as well, although in that case the map takes a set of matrices to the same set of matrices. The identity map is clearly a homeomorphism. Thus the facts about the space $E_{n}$ of cactus networks as shown in [33] and [17] are also true of the image of $\mathbf{\xi}$ in the compactified space of circular split systems. It is topologically a ball, and isomorphic to a certain subspace of the nonnegative Grassmannian. The map $\mathbf{\xi}$ often takes several cells of $E_{n}$ and embeds those cells inside of a single cell of the space of cactus circular split systems. However, the map does respect the cell structure in that, for cells, $x\leq y$ in the domain iff $\mathbf{\xi}(x)\leq\mathbf{\xi}(y)$ in the range. Thus it induces a new regular CW-complex structure on the circular electrical networks. Those cells in $E_{n}$ are labeled by their Kron reductions. Careful study of these cells in $E_{n}$ is important future work.
In this paper we first introduce cactus split systems and extend the maps to the domain of electrical cactus networks. Then we expand to global spaces: while traditionally the objects have a given cyclic order of $[n]$ we instead allow any cyclic order, and identify appropriate equivalent networks. That is how classical phylogentic spaces are studied, both the space of trees $BHV_{n}$ and the space of split networks (with all trivial splits) $CSN_{n}.$ The combinatorics become much more complex in the global case. With the equivalence in place even the basic enumeration of the objects becomes difficult. That is not surprising when we see how they relate to famous hard problems, specifically facets of the Symmetric Travelling Salesman polytope. Table 2 lists the total numbers of cells in our new global spaces, both compact and non-compact. The last column there lists the number of cells in the image of the global cactus networks, under our map $\mathbf{\sigma}$ that takes a cactus network to displayed splits on the parts of the partition into its connected components.
1 1 1 1 1 1
2 2 3 2 3 3
3 8 15 8 15 15
4 70 133 112 169 124
5 1466 2397 6976 ? ?
OEIS [?] [?] [?] [?] [?]
Table 2. Global spaces: numbers of cells.
@@ 1.2. Results
Some resulting numbers are shown in Table 1. The corollary follows via Lagrange inversion from the recursive functional equation of species composition, $A=B\circ(X\cdot A)$, as shown in Section 5. The general principle here is that the compactification $A$ of a space of networks is related to the original space of ordinary networks $B$ by that functional equation. Thus this corollary also applies to the counting of equivalence classes of electrical cactus networks: the number $|\mathbf{\Omega}_{n}|$ of cells for cactus networks can be found by:
@@
where $|\mathbf{\Omega}_{n}|$ is the number of distinct classes of non-compact electrical networks. For instance, (5(52)+10(8)2+10(2)3+10(4)+5(1))/5=(2(4)1)!!.55210821023104515double-factorial241(5(52)+10(8)2+10(2)3+10(4)+5(1))/5=(2(4)-1)!!.( 5 ( 52 ) + 10 ( 8 ) 2 + 10 ( 2 ) 3 + 10 ( 4 ) + 5 ( 1 ) ) / 5 = ( 2 ( 4 ) - 1 ) !! . The double factorial formula was shown in [33], but the larger formula is new. [^1] Our formula is implied by an inverse formula given by Paul D. Hanna in entry [A111088] of [37].
@@
For a planar cactus network $N$, the map $\mathbf{\xi}$ coincides with both the Kalmanson and the induced split systems of the planar dual. That is, $\mathbf{\xi}(N)=\mathbf{\rho}(N^{*})=\mathbf{\sigma}(N^{*}).$ Respectively, we have $\mathbf{\rho}(N)=\mathbf{\sigma}(N)=\mathbf{\xi}^{\prime}(N^{*}).$
Theorem 1 is a summary of Theorems 8 and 13 in Section 4, as extended to the compactified case via Theorem 16. Indeed, we show that the split systems in the image of this multi-named map have a particular form, known as faithful [14] — those systems which are made up of splits displayed by some network. These faithful systems are an important subcomplex of the space of circular systems. Theorem 1 has an immediate application for counting these faithful split systems. If the original network has multiple components, the faithful (induced) split system will be a cactus network, and vice versa. Thus counting the induced systems is more difficult. On the other hand, the graphical system has the same number of components and cactus form as the original system. Thus we can use our recursive counting principle to relate numbers of split systems for these faithful forms in the compact and non-compact cases, in Corollary 23 by the now recognizable formula:
$|\mathbf{\xi}(\mathbf{\Omega}_{n})|=|\mathbf{\sigma}(\mathbf{\Omega}_{n})|=\frac{1}{n+1}\left(\sum_{j_{0}+\dots+j_{n}=n}\left(\prod_{i=0}^{n}|\mathbf{\xi}(\mathbf{\Omega}_{j_{i}})|\right)\right).$
Section 5 proves enumeration results (summarized by Table 1, page 1) for our new bands in the spectrum of spaces (showcased by Figure 7, page 7). We define a class of plabic tilings for sets of polygons and then we finish with the following (expanding Corollary 23) :
@@
Finally, in Section 6 we introduce the globalization of the classical spaces and begin their study. We point out some initial results, such as the fact that the global spaces allow all cyclic orders, so their compactifications allow all partitions rather than just non-crossing partitions of $[n].$ Specifically, this means that the 0-cells of the CW-complex in each global case will be counted by the Bell numbers; the 0-cells in the non-global case are enumerated by the Catalan numbers. Section 7 collects some corollaries that relate our theorems to the global cases. We count the number of consistent cyclic orders for a network and describe obstructions to a given order being circular.
@@ 2.1. Circular Electrical Networks
Physically, a general electrical network $N$ is made of conducting Ohmic wires with $n$ exposed terminals that can be tested in order: We apply unit voltage to each of the terminals in turn while grounding all the remaining $n-1$ terminals. This defines the response matrix $M(N)$, recording the current at terminal $j$ of $N$ when the voltage is applied to terminal $i$ as entry $M_{ij}$. Indeed, two circular electrical networks are electrically equivalent if they have the same response matrix. An equivalence class is planar if any representative is planar.
@@
The Kron reduction $K(N)$ is an invariant of the equivalence class of $N$. Recall that equivalent networks are defined by having the same response matrix. For planar networks, planar representations of two in the same class are related by a sequence of moves including replacing a series of edges, replacing parallel edges, deleting superfluous edges, and the $Y-\Delta$ move. Those moves and several other invariants are discussed in [7]. The latter include the set of connections across the network, and the associated set of positive circular minors calculated from the response matrix—all of which are nonnegative for planar networks. Another invariant is the sub-collection of networks of a given equivalence class with the minimal number of edges, called the critical networks. We won’t need most of these invariants for this paper, but they are implicitly counted when we count the cells of the CW-complexes. Each cell is made of response matrices that all share the values of these (unweighted) electrical invariants. The weighted split system associated to a network is an electrical invariant as well, as shown in [13]. The unweighted set of splits is also an invariant of the electrical equivalence class, but like the unweighted Kron reduction it is strictly weaker than the set of connections or the set of unweighted critical networks.
@@
More practically, we could use an ohmmeter to test the resistance (impedance) between pairs of our terminals. We record these results as the resistance matrix $W(N)$, where $W_{ij}$ is the effective resistance between $i$ and $j.$ The entries $W_{ij}$ are a metric on the terminal nodes, as shown in [32].
Figure 3. (a) A nonplanar (with respect to the clockwise order) circular electrical network $N$, (b) its Kron reduction $K(N)$, and graphical split system $\mathbf{\xi}(N)$ in (c) polygonal and (d) split network formats.
Figure 3(a) shows a nonplanar case. Here are $M(N)$ and $W(N)$ for Figure 3(a):
@@
The space $\mathbf{\Omega}_{n}$ has dimension (${n\choose 2}.$ Each element in a cell of dimension $k$ in $\mathbf{\Omega}_{n}$ can be minimally represented by choosing positive conductance weights for a (non-unique) planar critical (or reduced) network with $k$ edges. A cell $x$ is contained in another cell $y$ when the network corresponding to $x$ can be found by either deleting or contracting edges of $y$. The left side of Figure 4 shows critical networks for each of the 52 cells of $\mathbf{\Omega}_{4}$. Each column corresponds to cells of $\mathbf{\Omega}_{4}$ of a fixed dimension with $f$-vector $(1,6,14,16,10,4,1)$.

### v4

# Compactifications of phylogenetic systems and species of electrical networks

@@ Abstract.
We describe new spaces and maps. Our graphical map is a visual and numerical correspondence between spaces of circular electrical networks and circular planar split systems. When restricted to the planar circular electrical case, this graphical map finds the split system uniquely associated with the Kalmanson resistance distance of the dual network, matching the induced split system familiar from phylogenetics. This correspondence is extended to compactifications of the respective spaces, taking cactus networks to the cactus split systems defined herein. The graphical map preserves both network components and cactus structure, allowing an elegant enumeration of induced phylogenetic split systems via combinatorial species. We introduce the global spaces of circular planar electrical networks and circular split systems. These new spaces are also CW complexes, but the 0-cells of each are counted by the Bell numbers as opposed to the Catalan numbers. As species, the two sorts of global cacti are seen to be compositions in complementary ways.
@@ 1.1. Motivation
A favorite tool of phylogenetics researchers is the split system, which is a combinatorial structure associated to a metric on a collection of biological species or taxa. The split system reveals the large-scale structure of the phylogenetic tree or network. Recently it was demonstrated that the split system is also guaranteed to exist in the case of a circular planar electrical resistor network [14]. Specifically, finding the split system is a consequence of a more basic result in that paper: the resistance metric for a circular planar electrical network is Kalmanson. This highlights a mathematical analogy between two problems: reconstructing a circuit in a ‘black box’ from surface measurements, and rebuilding a genetic history from existing population data.
@@
Figure 1. Example labels for the cells in the spaces we consider, for $n=4$: (a) circular planar electrical networks, (b) cactus networks (c) circular split systems and (d) circular cactus split systems . The total numbers of these cells are listed in Table 1, row 4, and the entire sets for (a) and (c) are shown in Figure 5.
@@
In this paper we carry the analogy further. A printed electrical circuit is made of conductors in a network, with terminal nodes on the exterior available for connections. Some of those exterior nodes might be duplicates, short-circuited together so that they are interchangeable in terms of their use. Similarly, monozygotic twins can be seen as genetically the same for the purpose of reconstructing a family tree from the DNA samples of living individuals. On a larger scale, basic DNA sequences of extant individuals from a diverse ecosystem can be equated when they come from the same biological species. Allowing nodes, either terminals or taxa, to be identical in the limit of being infinitely close to each other corresponds to taking an unbounded space and compactifying it. Here we define and relate the compactification of phylogenetic split systems (from the DNA examples) to the well-studied compactification of circular planar electrical network space.
@@
1 1 1 1 1 1 1
2 2 3 2 3 2 3
3 8 15 8 15 8 15
4 52 105 64 117 49 102
5 464 945 1024 1565 373 839
OEIS [A111088] [A001147] [A006125] [A136654] [?] [?]
Table 1. Spaces: numbers of cells in each CW complex. Left to right these are circular planar electrical networks, cactus planar electrical networks, circular split systems, and circular cactus split systems. The last two columns count the cells in the range of our maps from electrical networks to split systems. Examples are in Figure 1.
1.2. Background
In our recent papers [14, 15], we demonstrated an injective mapping from electrical equivalence classes of circular planar electrical networks to circular split systems. Every planar response matrix is a Kalmanson metric, and thus associated to a unique weighted circular split system. To see if a given circular split system is the image of an electrical network is more complicated. For a given Kalmanson metric one can first reorder the matrix to match any cyclic order respected by its circular split system. Then there is a unique associated response matrix, which can be checked for nonnegative circular minors to decide if it is a planar electrical network. If so, then we show in [14] that the circular planar electrical network has the same overall tree structure as the split system network of the resistance matrix. Then the algorithm from Chapter 9 of [8] can reconstruct a critical subnetwork for any part not made of bridges, using the strand diagram. In [25] the authors show how to check for planarity directly using the (reordered) resistance metric. This is done via the Lam embedding into the nonnegative Grassmannian, but with a variation that takes the resistance metric as input. Then in [25] the authors show an alternative reconstruction of the electrical network again using the resistance metric directly to find the strand matching.
Figure 2. Domains and ranges for the maps in this paper.
1.3. Overview
From a geometric combinatorial perspective, our interest lies in comparing the spaces of circular planar electrical networks and circular split systems as CW-complexes with analogous cell-decompositions. We will review the definitions in the following sections, but for now we show some examples of each, to introduce our notation. $\mathbf{\Omega}_{n}$ is the space of circular planar electrical networks with $n$ exterior nodes, in notation from [8]. $\mathbf{\Omega}_{n}$ has its poset of cells denoted $EP_{n}$ in [1]. The cells correspond to equivalence classes of critical networks such as (a) in Figure 1. The compactification of $\mathbf{\Omega}_{n}$ is the space of cactus networks, denoted $E_{n}$ in [35]. We denote by $\mathbf{\Omega}_{n}$ the poset of cells of $E_{n},$ with an example in (b) of Figure 1.
The space of split systems on the set $[n]$ is denoted $\mathbf{\Psi}_{n}.$ Subfigure 1 (c) shows an example of a split system in $\mathbf{\Psi}_{4}.$ The compactification of $\mathbf{\Psi}_{n}$ is defined in this paper and is denoted $\mathbf{\Psi}_{n}$, with an example in (d). In the interest of readability, we often use one name for both the CW-complex and the poset of cells: we will refer to $|\mathbf{\Omega}_{n}|$ as the number of cells in $\mathbf{\Omega}_{n}$ and to $|\mathbf{\Psi}_{n}|$ as the number of cells in $\mathbf{\Psi}_{n}.$ Similarly, we will use $\mathbf{\Omega}_{n}$ interchangeably with $E_{n}.$
Further, we want to consider the sub-complexes of split systems that are the range of the graphical map $\mathbf{\xi}$, which takes circular planar electrical cactus networks to circular cactus split systems. In Figure 1 (c) is the image of (a) under $\mathbf{\xi}$, and (d) is the image of (b). There are fewer cells in that range than split systems in general: the implied mapping on cells is not onto. Neither is it an injection of cells. However the graphical map does preserve the connected components and cactus structures, and this provides us a way to count cells in the range either in either case, whether we want just the ordinary split systems or their compactifications as well. Table 1 shows the results of counting the cells of all dimension in the various complexes.
The map $\mathbf{\xi}$ is quite simple in its operation on matrices: it is the identity map (on the independent portion of the matrix.) The identity map is clearly a homeomorphism. Thus the facts about the space $E_{n}$ of cactus networks as shown in [35] and [18] are also true of the image of $\mathbf{\xi}$ in the compactified space of circular split systems. It is topologically a ball, and isomorphic to a certain subspace of the nonnegative Grassmannian. The map $\mathbf{\xi}$ often takes several cells of $E_{n}$ (elements of $\mathbf{\Omega}_{n})$ and embeds those cells inside of a single cell of the space of circular cactus split systems, $\mathbf{\Psi}_{n}$. However, the map does respect the cell structure in that, for cells, $x\leq y$ in the domain iff $\mathbf{\xi}(x)\leq\mathbf{\xi}(y)$ in the range. Thus it induces a new regular CW-complex structure on the circular electrical networks. Those new cells in $E_{n}$ are labeled by their Kron reductions, or by their image in $\mathbf{\Psi}_{n}$ which we will designate by plabic tilings. Careful study of these cells in $E_{n}$ is important future work.
In this paper we first introduce cactus split systems and extend the maps to the domain of electrical cactus networks. Then we expand to global spaces: while traditionally the objects have a given cyclic order of $[n]$ we instead allow any cyclic order, and identify appropriate equivalent networks. That is how classical phylogentic spaces are studied, both the space of trees $BHV_{n}$ and the space of split networks (with all trivial splits) $CSN_{n}.$ The combinatorics and topology become much more complex in the global case. With the global equivalence in place even the basic enumeration of the objects becomes difficult. That is not surprising when we see how they relate to famous hard problems, specifically facets of the Symmetric Travelling Salesman polytope. Table 2 lists the total numbers of cells in our new global spaces, both compact and non-compact. The last column there lists the number of cells in the image of the global cactus networks, under our map $\mathbf{\sigma}$ that takes a cactus network to displayed splits on the parts of the partition into its connected components.
1 1 1 1 1 1
2 2 3 2 3 3
3 8 15 8 15 15
4 70 133 112 169 124
5 1466 2397 6976 7857 ?
OEIS [?] [?] [?] [?] [?]
Table 2. Global spaces: numbers of cells.
@@ 1.4. Results
Some resulting numbers are shown in Table 1. The formula in Corollary 18 arises via Lagrange inversion from a recursive functional equation of combinatorial species composition. We use the same notation to refer to species and generating functions as we use for spaces and CW complexes. Thus for circular cactus split systems the species equation is $\mathbf{\Psi}=\mathbf{\Psi}\circ(X\cdot\mathbf{\Psi})$, as shown in Section 5. The general principle we notice is that often the compactification of a space of networks is related to the original space of ordinary networks by that same functional equation. This corollary also applies to the counting of equivalence classes of electrical cactus networks. Since as species we have $\mathbf{\Omega}=\mathbf{\Omega}\circ(X\cdot\mathbf{\Omega})$ the number $|\mathbf{\Omega}_{n}|$ of cells for cactus networks can be found by:
@@
where $|\mathbf{\Omega}_{n}|$ is the number of distinct classes of non-compact electrical networks. For instance, (5(52)+10(8)2+10(2)3+10(4)+5(1))/5=(2(4)1)!!.55210821023104515double-factorial241(5(52)+10(8)2+10(2)3+10(4)+5(1))/5=(2(4)-1)!!.( 5 ( 52 ) + 10 ( 8 ) 2 + 10 ( 2 ) 3 + 10 ( 4 ) + 5 ( 1 ) ) / 5 = ( 2 ( 4 ) - 1 ) !! . The double factorial formula was shown in [35], but the larger formula is new. [^1] Our formula for $(2n-1)!!$! here is related to an inverse formula given by Paul D. Hanna in entry [A111088] of [38].
@@
For a planar cactus network $N$, the graphical system $\mathbf{\xi}(N)$ coincides with both the Kalmanson and the induced split systems of the planar dual $N^{*}$. That is, $\mathbf{\xi}(N)=\mathbf{\rho}(N^{*})=\mathbf{\sigma}(N^{*}).$ Respectively, we have $\mathbf{\rho}(N)=\mathbf{\sigma}(N)=\mathbf{\xi}^{\prime}(N^{*}).$
We show the domains and codomains of the maps in Figure 2. Theorem 1 is a summary of Theorems 8 and 13 in Section 4, as extended to the compactified case via Theorem 16. Indeed, we show that the split systems in the image of this multi-named map have a particular form, known as faithful [15] — those split systems which are made up of splits displayed by some circular planar network, phylogenetic or electrical. These faithful systems are an important subcomplex of the space of circular split systems. Since taking duals of a circular planar split network is a bijective operation, Theorem 1 has an immediate application for counting these faithful split systems. We note that if an electrical cactus network $N$ has multiple connected components, then the faithful (induced) split system $\mathbf{\sigma}(N)$ will have cactus bulbs corresponding to connected components of $N$ and vice versa. This complicates directly counting the induced split systems. On the other hand, our graphical split system $\mathbf{\xi}(N)$ has the same number of components and the same cactus form as the network $N$. Thus we can use our recursive counting principle to relate numbers of split systems for these faithful forms in the compact and non-compact cases, in Corollary 23 by the now recognizable formula:
Section 5 proves enumeration results (summarized by Table 1, page 1) for our spectrum of spaces (showcased by Figure 2, page 2). In order to recognize the form of faithful split systems we define a new class of plabic tilings for sets of polygons. Then we finish with the following (expanding Corollary 23) :
@@
In Section 6 we introduce the globalization of the classical spaces and begin their study. We point out some initial results, such as the fact that the global spaces allow all cyclic orders, so their compactifications allow all partitions rather than just non-crossing partitions of $[n].$ In Theorem 29 we see that the global cactus split systems can be described as a composition of species: $\mathbf{\Psi}^{global}=E_{+}\circ\mathbf{\Psi}^{global}$ where $E_{+}$ is the species of non-empty sets. That is, a global cactus split system is formed by partitioning the set $[n]$ and then making a global ordinary split system on each part. As exponential generating functions we have:
${\mathbf{\Psi}^{global}(x)=e^{\mathbf{\Psi}^{global}(x)}-1.}$
In contrast, as a species the global cactus electrical networks are described by the opposite composition: $\mathbf{\Omega}^{global}=\mathbf{\Omega}^{global}\circ E_{+}$. As exponential generating functions we have:
$\mathbf{\Omega}^{global}(x)=\mathbf{\Omega}^{global}(e^{x}-1).$
Section 7 collects some corollaries that relate our theorems to the practical questions (1)-(4), via obstructions to circular planarity. Using the plabic tilings we count the number of consistent cyclic orders for a planar network and connect that collection to a face of the Symmetric Traveling Salesman polytope.
@@ 2.1. Circular Electrical Networks
Physically, a general electrical network $N$ is made of conducting Ohmic wires with $n$ exposed terminals that can be tested in order: We apply unit voltage to each of the terminals in turn while grounding all the remaining $n-1$ terminals. This defines the response matrix $M(N)$: entry $M_{ij}$ is the current at terminal $j$ of $N$ when the unit voltage is applied to terminal $i$.
@@
Two circular electrical networks are electrically equivalent if they have the same response matrix. An equivalence class is planar if any representative is planar.
@@
The Kron reduction $K(N)$ is an invariant of the electrical equivalence class of $N$. The weighted split system associated to a network is an electrical invariant as well, as shown in [14].
For circular planar networks, planar representations of two networks in the same class are related by a sequence of moves selected from: 1) replacing a series of edges, 2) replacing parallel edges, 3) deleting superfluous edges, and 4) the $Y-\Delta$ move. Those moves are pictured in [8]. Also discussed in that source are several other weaker invariants of electrical equivalence, including the set of connections across the network, and the associated set of positive circular minors calculated from the response matrix. All the circular minors must be nonnegative for planar networks, but which of them are positive or zero depends on the specific network. Another invariant is the sub-collection of networks of a given equivalence class with the minimal number of edges, called the critical networks. We do not directly use the definitions of connections and circular minors in this paper, but they are implicitly considered when we enumerate the cells of $\mathbf{\Omega}_{n}$. Each cell is made of response matrices that share values of these (unweighted) electrical invariants: the collections of connections, positive minors, and critical networks.
The unweighted set of splits is also an invariant of the electrical equivalence class, as shown in [14], but like the unweighted Kron reduction, it is strictly weaker than the set of connections or the set of unweighted critical networks.
@@
As an alternative measurement we could use an ohmmeter to test the resistance (impedance) between pairs of our terminals. We record these results as the resistance matrix $W(N)$, where $W_{ij}$ is the effective resistance between $i$ and $j.$ The entries $W_{ij}$ are a metric on the terminal nodes, as shown in [33].
Figure 4(a) shows a non-planar case.
Figure 4. (a) Non-planar (with respect to the clockwise order) circular electrical network $N$.
Then we show $M(N)$ and $W(N)$ for Figure 4(a):
@@
The space $\mathbf{\Omega}_{n}$ has dimension (${n\choose 2}.$ Each element in a cell of dimension $k$ in $\mathbf{\Omega}_{n}$ can be minimally represented by choosing positive conductance weights for a (non-unique) planar critical (or reduced) network with $k$ edges. A cell $x$ is contained in another cell $y$ when the network corresponding to $x$ can be found by either deleting or contracting edges of $y$. The left side of Figure 5 shows critical networks for each of the 52 cells of $\mathbf{\Omega}_{4}$. Each column corresponds to cells of $\mathbf{\Omega}_{4}$ of a fixed dimension, 0–6, with $f$-vector $(1,6,14,16,10,4,1)$. The $f$-vector of $\mathbf{\Omega}_{5}$ is $(1,10,40,85,110,97,65,35,15,5,1).$